40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 27 20:47:00 2007, tobin, Update, PSL, FSS fss-tf0001.pdf
    Reply  Wed Nov 28 04:21:46 2007, rana, Update, PSL, FSS 
Message ID: 127     Entry time: Tue Nov 27 20:47:00 2007     Reply to this: 128
Author: tobin 
Type: Update 
Category: PSL 
Subject: FSS 
Rana, Tobin

We looked at the RF PD signal to the FSS (siphoning off a signal via a minicircuits directional coupler) and also took an open loop transfer function of the FSS. In the transfer function we saw the step at 100 kHz (mentioned by Rob) as well as some peculiar behavior at high frequency. The high frequency behavior (with a coupling of ~ -20 dB) turns out to be bogus, as it is still present even with the beam blocked. Rearranging the cabling had no effect; the cause is apparently inside the FSS. The step at 100 kHz turns out to be a saturation effect, as it moved as we lowered the signal amplitude, disappearing as we approached -60 dBm. (Above the step, the measurement data is valid; below, bogus.)

Transfer functions will be attached to this entry.

Some things to check tomorrow: the RF signal to the PC, RF AM generation by the PC, LO drive level into the FSS, RF reflection from the PC, efficiency of FSS optical path, quality of RF cabling.
Attachment 1: fss-tf0001.pdf  761 kB  Uploaded Wed Nov 28 15:34:07 2007  | Hide | Hide all
fss-tf0001.pdf fss-tf0001.pdf
ELOG V3.1.3-